how much money should be invested in an account that earns 9% interest, compounded monthly, if you want to have 5000000 after 20 years
Answer by mananth(16946)

Principal P = x
Amount= 5000000
years=n 20
compounded 12 times a year (t)
Rate = 9.00 0.750
Amount = P*((n+r)/n)^nt

5000000 = x *( 1 + r/n )^ n


5000000 =P(1+0.090/12)^ 12* 20

5000000 = P( 1.00750 )^ 240

5000000 /( 1.01 )^ 240 = P

P= 832064.22